# Correlated Growth of Causal Networks

### v1

@@ Abstract
averaging Eq. (6) over the latent variable $\lambda$. Figure 2 plots the normalized joint degree distribution $R(k,k^{\prime})\mathrel{\mathop{:}}=\frac{P(k,k^{\prime})}{P_{rand}(k,k^{\prime})}$, i.e., $P(k,k^{\prime})$ compared to the randomized one [28]. $R>1,=1,<1$ indicates correlated, uncorrelated, and anticorrelated degrees, respectively. Our results show that the theoretical prediction successfully captures the assortative nature of the citation graph, aligning well with empirical measurements.
Figure 3: K-nearest-neighbor vs degree $k^{\prime}$. This figure shows the empirical data (scatters) and modeling predictions (solid lines) of $k_{nn}(k_{0})$ across four real networks.
To compare our theory with empirical data more quantitatively, we calculate the $k$-nearest-neighbor $k_{nn}(k^{\prime})\mathrel{\mathop{:}}=\sum_{k}kP(k|k^{\prime})=\sum_{k}\frac{kP(k,k^{\prime})}{P(k^{\prime})}$, which measures the dependency of a node’s degree on those of its neighbors. The scatter plot in Fig. 3 shows that $k_{nn}$ increases with $k^{\prime}$ and converges to finite values for large $k^{\prime}$, suggesting a clear positive correlation. Our theoretical prediction, using Eq. (9), depicted by the solid curves, shows good agreement for all four networks. Interestingly, the saturation for large $k^{\prime}$ is non-trivial. Our theory indicates this is not a finite-size effect but is instead associated with both dynamical processes and the fitness distribution. Indeed, our theory shows
$k_{nn}(k^{\prime}\to\infty)\sim\frac{\langle k\rangle\langle\lambda e^{-k_{0}\lambda}\rangle_{\text{lat}}\langle\phi(t)\rangle_{t}}{r},$ (10)
@@
Table 1: Degree correlation coefficients
Subjects Empirical Data Theory
@@
Conclusion. Our study demonstrates that topological correlations in causal networks inherently emerge from the underlying dynamics, driven by the intrinsic causal and dynamical correlations associated with network growth. We validate our theoretical framework using citation graphs as a representative case. In this context, our approach addresses notable limitations of existing models [12, 11] by reducing the number of free parameters from many to a few, thereby underscoring the efficacy of our methodology. Moreover, by integrating appropriate causal correlation kernels, our framework is readily applicable to a diverse array of causal networks, including those in social media, biological evolution, and economic growth.
Several questions remain open, with primary challenges for future research including the incorporation of time-varying state distributions, particularly in causal networks exhibiting non-stationary growth. Additionally, exploring topological correlations that extend beyond the nearest-neighbor level constitutes another important direction for future investigation. Addressing these challenges will enhance our understanding of how causal correlations influence network dynamics across various domains. Nonetheless, our findings provide a promising pathway for advancing the comprehension of correlations in causal networks, with the potential to significantly impact fields such as physics, economics, social sciences, and evolutionary biology.

### v2

@@ Abstract
Moreover, the joint degree distribution $P(k,k^{\prime})$ enables the analytical computation of the assortativity coefficient $r_{corr}\mathrel{\mathop{:}}=\frac{Cov(k,k^{\prime})}{\sqrt{Var(k)Var(k^{\prime})}}$ [35]. Using the law of total covariance and the marginal independence of $k$ and $k^{\prime}$, Eq. (7) naturally leads to (see SM Section 1.6)
$r_{corr}=\frac{Cov\left(\overline{k}(\lambda,\tau),\overline{k}(\lambda^{\prime},\tau^{\prime})\right)}{\sqrt{Var\left(\overline{k}(\lambda,\tau)\right)Var\left(\overline{k}(\lambda^{\prime},\tau^{\prime})\right)}},$ (10)
as the correlation coefficient over $\mathbf{\Psi}$. Table (1) summarizes the assortativity $r_{corr}$ for both empirical results and theoretical predictions, demonstrating good agreement.
Table 1: Assortativity coefficients
Subjects Experiment Theory
@@
Discussion. Our study demonstrates that topological correlations in causal networks naturally originate from fundamental causal and dynamical correlations during network growth. Analytical results explicitly reveal that these correlations induce marginal dependencies, which in turn generate observable degree correlations. We validate our theoretical framework by applying it to empirical citation graphs, which serve as representative real-world examples.
Compared to existing fitness-based models, which assign a separate fitness value to each vertex and require $O(N)$ parameters for a network with $N$ vertices, these models often suffer from overfitting, thereby limiting predictive power. In contrast, our model introduces causal dependencies between fitness values, encoded in a causal kernel, which reduces the number of free parameters from $O(N)$ to $O(1)$. This structured approach mitigates overfitting while improving both explanatory and predictive capabilities. In this way, our method addresses notable limitations of existing models [12, 11], demonstrating that causal dependencies provide a more compact and generalizable framework for modeling real-world networks.
While our theory successfully captures key causal dependencies, several open questions remain. One major challenge is the incorporation of time-varying state distributions, particularly in causal networks exhibiting non-stationary growth. If Equation (4) admits no solution, stationarity breaks down. This can occur when the causal kernel has sufficiently broad tails. Consequently, the network undergoes super-exponential growth, violating TTI. This scenario is closely related to the “winner-takes-all” phase, where a small fraction of vertices accumulate a disproportionate share of influence, analogous to Bose-Einstein condensation [25]. Since this behavior does not arise in the empirical data we analyze, we defer its detailed investigation to future work. Another important direction is to explore topological correlations beyond the nearest-neighbor level. Higher-order structural dependencies may reveal additional causal mechanisms governing network formation, further enriching our understanding of dynamical correlations in complex systems.
Despite these open challenges, our findings provide a robust foundation for understanding correlations in causal networks. By incorporating causal correlation kernels, our framework applies to a wide range of systems, including social media, biological evolution, and economic growth. This approach offers new insights into the interplay between causal structure and network topology, with broad implications across physics, economics, social sciences, and biology.